OpenStudy (anonymous):

for the last part of my comp assignment, we have to come up with an interesting sequence that would be cool to visualize on a graph...You can use numbers, "n"'s (be the natural numbers, changing as n gets bigger/smaller), fibonacci sequences in it.... any ideas??

OpenStudy (anonymous):

try the sum of the series with general term 1/r, where r varies from 1 to n.let s be the sum. s=1/1+1/2+1/3+.....+1/n. plot s against n. this is a oscillating type of graph.

OpenStudy (mrdoe):

do a power series for a function, like \[\sin x = \sum_{n=0}^{\infty} \frac{ (-1)^n x ^{2n} }{ 2n! }\] That's sure to impress your prof